Report for July 2018

Report Date
July 2018

Section One: Programme Commentary, RAG and Resources Update

Programme Commentary

There is a total of four projects still open at the end of the 17-18 session, and these will all continue into 18-19 to varying dates.

The first - COM036 - now looks it may have found a solution to the performance issues that had hindered progress over the last few months; this is due to some in-depth analysis of our infrastructure by ITI and Production Management. We are expecting to see some changes made in August that move things along. There has also been progress made on the remaining steps of this project and these have been planned out, with an eventual delivery in December '18.

COM041 is now expected to deliver in November '18, with the remaining milestones moved out because of the issues in scheduling migration work round annual leave; staff availability, and the freeze change. The remaining projects are looking to complete in the shorter term - COM042 has had a small delay in being signed off, and COM040 looks like it will deliver an upgrade to JIRA by the end of August. It should be pointed out, however, that there is still a question around the Oracle version for this.

 

RAG Status and Forward Look

 

Resources

 

 Section Two: Projects Not Started

Project Info     Project Manager's Commentary
N/A

 

 

Section Three: Projects in Progress

 

Project Info Project Manager's Commentary

COM036 BI - Business Objects Upgrade

Compliance

Milestones

05-Oct-2018 DEV build complete
19-Oct-2018 Test build complete
26-Oct-2018 Integration Testing sign off
21-Nov-2018 Acceptance Testing sign off
30-Nov-2018 Live build complete
11-Dec-2018 Deploy to LIVE
22-Jan-2019 Deployment Sign Off Review
29-Jan-2019 Closure

RAG Status 

AMBER

 

Report by Chris Konczak for July 2018

Achievements in Last Period

Good progress made by Production and ITI on the performance issues with the following tasks identified and to be started in July, with a completion in mid August:

  • ITI to replace all the old 3Com network switches - there are definite correlations between discards and errors on the switch uplinks that relate to latencies and errors seen in vCenter and this will affect any application making network connections
  • ITI to replace the older hosts in the cluster - this is planned work and is already underway

These are likely to have a positive impact on the BI performance. Once these are put in place the project team will retest and feedback on the results.

  • ODBC / JDBC connectors updated
  • Issues with Oracle Client and OCI components resolved
  • Milestones replaned and budget agreed for continuing the project in 2018/19

Issues

  • Project status now Amber as ITI and Production have identified the potential areas that could be causing the problems and have put in place a plan to resolve.
  • Funding to be resolved and approved for £6.5K to have ITelligence install the patch or service pack to allow content migration using the wizard. Programme Owner has indicated that they would be able to fund this. Without this, the current situation is that the content migration is carried out using the Promotion Manager which is untried.

Next Steps

  • ITI to replace all the old 3Com network switches
  • ITI to replace the older hosts in the cluster
  • Project team to test and feedback

 

COM038 SOA Dev Upgrade to 12c

Compliance

 

Milestones

Project is now closed

 

RAG Status

BLUE

Report by Adam Wadee for July 2018

Achievements

  • Remaining milestones signed off
  • Project closure completed

Issues

  • None

Next steps

  • None

 

COM040

JIRA Upgrade

Compliance

Milestones

14-Aug-2018 Build review Build
29-Aug-2018 Acceptance review Accept
11-Sep-2018 Deployment review Deliver
27-Sep-2018 Closure review Close

RAG Status

Green

Report by David Watters for July 2018

Achievements

  • Project brief completed and signed off
  • DEV and TEST refreshed from current LIVE
  • DEV upgraded and tested
  • TEST upgraded and testing commenceds

Issues

  • Question around version of Oracle to be used - still to be decided

Next steps

  • Sign-off DEV & TEST upgrades
  • Decide on database version
  • Yes/No meeting re deployment to LIVE

COM041 Decommission legacy Windows hosting

Compliance

 

Milestones

Target Date Title
23-Nov-2018 Migration complete
07-Dec-2018 Closure complete

 

RAG Status

GREEN

 Report  by Tim Gray for July 2018

 

Achievements in Last Period

  • Good progress made on easier sites.
  • Committee Intranet site majority are already migrated, these need to be cross-checked.

Issues

  • Due to the time taken so far to migrate sites, annual leave and the change control period, this project will now move into November --> December closure

Next steps

  • Continue with site migrations and checking.

 

 

 

COM042 GDPR Preparation for Applications

Compliance

Milestones

25-May-2018

Acceptance of Action Plan and Final Report

10/8/18

Project Closure

 

RAG Status

AMBER

 

Report by Emma Mcnab for July 2018

 

Report for July 2018

Report Date
July 2018

Achievements

  • Action plan progressed

Issues 

  • Milestones for closure have slipped -  PICCL required to outline how closure will be reached before end of year.

Next Steps

  • Closure report / project close

Amber: Change - Closure milestone has slipped due to availability of Project Manager / Sign of from Sponsor.  Project Manager  undertaking review of milestones however project is expected to close w/c 6/08/18.

 

 

Section Four: Closed or Withdrawn Projects

Project Info Project Managers Commentary

COM034

Annual Wiki Upgrade

Compliance

Report by Chris Konczak

This project was closed on 08/08/17  (WIS Notes 09/09/17)

Please see the Completion Report for further information

COM035 

JIRA Upgrade 2017

Compliance

Report by Tim Gray

This project was closed on 14/08/17  (WIS Notes 11/08/17)

Please see the Completion Report for further information

COM032

Rollout Academic Timetable Integration with Office 365 Calendar

Discretionary

Report by Chris Copner

This project was closed on 18/08/17  (WIS Notes 18/08/17)

Please see the Completion Report for further information

COM031

Serengeti Document Management Server Upgrade

Compliance

Report by Morna Findlay

This project was closed on (WIS Notes 23/03/18)

Please see the Completion Report for further information.

COM037

Grouper Upgrade

Compliance

Report by David Watters

Project was closed on 22/06/18 (WIS notes 22/06/18)

Please see the completion report for further information

COM039

Wiki Upgrade

Compliance

Report by Chris Konczak

This project was closed on 20/07/18  (WIS Notes 20/07/18)

Please see the Completion Report for further information

 

Section Five: Actions/Notes from monthly meeting held on 15/05/2018

Project Action / Notes Responsibility    Status
COM036

The migration tool is not fit for purpose. SC to await feedback from vendor today if not successful then escalate through AC

PJ to speak to Mark McG re figures on the technical issues. PJ suggested stress test on LIVE with 5/6 users - SC to speak to AMcF to organise.

CK to progress weekend work - has everyone responded? Have ITI confirmed their availability for DNS change? Project team to discuss if weekend work is really necessary?

SC/AC

PJ/SC/AMcF

TG/CK

 
COM038

Project should aim to complete as much infrastructure work this year as possible to utilise the available resources. 

AW

 
COM039 License issue to be clarified - TG follow up with CK/SK TG/CK  
COM040 JIRA Upgrade project to be moved to start in June. Forward Look to be updated and PRICCL raised. DW  

 

AttachmentSize
Image icon Reources for July 2018189.33 KB
Image icon Forward Look July 18117 KB